# Quality control in sublinear time: a case study via random graphs

### v1

@@ 1.1 Quality Control problems
For function $\rho:\mathcal{X}\to\mathbb{R}_{\geq 0}$, small constant $\varepsilon>0$, and a distribution $\mathcal{D}:\mathcal{X}\to\mathbb{R}_{\geq 0}$ such that $\mathbb{P}_{x\sim\mathcal{D}}\left[\left|\rho(x)-1\right|\leq\varepsilon\right]\geq 9/10$, a $(\mathcal{D},\rho)$-quality control algorithm $A$ must satisfy the following:
@@
    $\mathbb{P}_{x\sim\mathcal{D};R}\left[A(x)=\textsc{Accept}\right]\geq 2/3$, where the probability is taken over distribution $\mathcal{D}$ and the randomness $R$ used by algorithm $A$.
@@ 1.3 Comparison to related works
Given this proximity to the notions of average case complexity, one may anticipate seeing instances of quality control problems arising naturally in the field. Indeed, one notable class of such problems is “refutations of random 3SAT” problems. Here a seminal definition of Feige [Fei02] requires that a refutation algorithm for 3SAT problem on distribution $\mathcal{D}$ should only accept unsatisfiable 3SAT inputs, while accepting most inputs drawn from $\mathcal{D}$. This corresponds exactly to a quality control problem with the parameter $\rho_{\mathrm{SAT}}(\phi)$ being $1$ if a formula $\phi$ is unsatisfiable and 0 otherwise. Additionally, the problem of certifying properties of a random object has appeared in various contexts throughout theoretical computer science. Notably, natural proofs [RR97] examine such a problem (within the context of ruling out avenues towards super-polynomial circuit lower bounds). Additionally, as noted by [Tre06], certifying properties of random instances is a challenge and a crucial step in explicit constructions from random objects. The problem of certifying properties of random objects in these various settings can be framed as a quality control problem.
@@ 2.3 Properties of Erdős–Rényi graphs
For $G\sim\mathcal{G}_{n,p}$, with probability at least $9/10$, the number of labeled induced copies of $H$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{v(H)}v(H)!\cdot p^{e(H)}(1-p)^{\binom{v(H)}{2}-e(H)}$.
@@
For $G\sim\mathcal{G}_{n,p}$ and $p=\Omega\left(n^{-2/(k-1)}\cdot\varepsilon^{-2}\right)$, the number of $k$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{k}k!\cdot p^{\binom{k}{2}}$ with high constant probability (say, $9/10$).
@@
For all $p=\Omega\left(\frac{1}{n}\right)$, there exists an $N$ such that for all $n\geq N$, $\mathbb{P}_{G\sim\mathcal{G}_{n,p}}\left(\alpha(G)\leq 2\frac{e(G)}{n-1}+1\right)\geq 9/10$.
@@
$\mathbb{P}_{G\sim\mathcal{G}_{n,p}}\left(\alpha(G)\leq 2np\right)\geq 4/5.$
@@
Let $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$ and $s\geq 1200C^{4}/p^{5k}$. Let $\delta=\frac{p^{2k}}{12C^{2}}$. With probability at least $9/10$, for $G\sim\mathcal{G}_{n,p}$, at least a $9/10$-fraction of subgraphs $S$ of size $s$ satisfy the following:
@@
Therefore, the expected number of $S\subseteq[n],|S|=s$ such that $G_{S}$ is not good is at most $\frac{1}{s}\cdot\binom{n}{s}\leq\frac{1}{100}\binom{n}{s}$ for $s\geq 100$.
For a graph $G$, let $t(G)$ be the number of $S\subseteq[n],|S|=s$ such that $G_{S}$ is not good. By Markov’s inequality,
$\mathbb{P}\left(t(G)\geq\frac{1}{10}\cdot\binom{n}{s}\right)\leq\frac{\mathbb{E}(t(G))}{\frac{1}{10}\cdot\binom{n}{s}}\leq\frac{\frac{1}{100}\cdot\binom{n}{s}}{\frac{1}{10}\cdot\binom{n}{s}}\leq\frac{1}{10}.$
Therefore, at least a $9/10$-fraction of subgraphs of size $s$ are good. ∎
@@
Let $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$, $s\geq 1200C^{4}/p^{5k}$, and $\delta=\frac{p^{2k}}{12C^{2}}$. Consider $G\sim\mathcal{G}_{n,p}$, and choose a random subgraph $G_{S}$ of $G$ on $s$ vertices. Then, with probability at least $4/5$, the subgraph on these $s$ vertices satisfies the following:
@@
This corollary follows by taking a union bound over the events in Section 2.3 that $G\sim\mathcal{G}_{n,p}$ does not satisfy the property that at least a $9/10$-fraction of size-$s$ subgraphs are good and that a non-good subgraph is chosen at random. ∎
@@ 3 Quality control of k-clique count
    When $p=\Omega(\log(n)/n)^{1/(4k-1)}$, the problem is additionally solvable in $O(p^{-c_{3}k})$ time.
@@ 3.1.2 Proof of Inductive Lemma given Concentration Lemma
Suppose that a graph $G$ satisfies the property that for all $\ell\in\{2,3,\dots,k\}$ the number of labeled $\ell$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{4(k+1)})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$ (for $p=\Omega\left(\varepsilon^{-2}n^{-2/(k-1)}\right)$). Then, for size $s_{\ell}$ multisets, for all $\ell\in[k-1]$, $G$ is $\left((\ell+1)\cdot\frac{\varepsilon}{4(k+1)},(\ell+1)^{2}\cdot\frac{\varepsilon^{2}}{(k+1)^{2}}\cdot\frac{p^{2\ell-2}}{2048},s_{\ell}\right)$-exponentially robustly quasirandom with respect to $\ell$-clique counts.
@@
Completeness. By Section 2.3, with high constant probability, for $G\sim\mathcal{G}_{n,p}$ with $p=\Omega\left(\varepsilon^{-2}n^{-2/(k-1)}\right)$, for all $\ell\leq k$ the number of labeled $\ell$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{4(k+1)})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$. In the following, suppose that the high-probability event that the clique counts are all in these typical ranges holds. Then, by Claim 3.1.3, for all $\ell\in[k]$, $G$ is $\left(\delta_{\ell},\nu_{\ell},s_{\ell}\right)$-exponentially robustly quasirandom with respect to $\ell$-clique counts, for
@@
By Section 3.1.2, from this exponential quasirandomness, we can conclude that, with probability at least $1-k\cdot 4^{k}\exp\left(-\nu_{k}s_{*}\right)\geq 9/10$, for all $\ell\in[k]$, the uniformly sampled multiset $S$ satisfies $C_{\ell}(S)\in(1\pm\frac{\varepsilon}{4})E_{s,\ell}$. Therefore, the algorithm accepts $G\sim\mathcal{G}_{n,p}$ with high constant probability.
@@ 3.3 Algorithmic complexity
Completeness. We first argue that, with high probability, $G\sim\mathcal{G}_{n,p}$ is accepted by Algorithm Clique-Quality-Efficient. Consider $G\sim\mathcal{G}_{n,p}$, and choose a random subgraph of $G$ on $s_{*}$ vertices. By Corollary 2.3, for $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$ and $s_{*}\geq 1200C^{4}/p^{5k}$, with probability at least $4/5$, the subgraph on these $s_{*}$ vertices satisfies the conditions checked for in Step 3 with high probability, as proven in Section 2.3. Therefore, Step 3 does not reject $G\sim\mathcal{G}_{n,p}$ with high constant probability.
@@ 4.1 Query complexity upper bound
Define $s$ as in Section 5. Then $s=\widetilde{O}\left(r_{k}(H)^{-1}\varepsilon^{-2}\right)=\widetilde{O}\left(p^{-2\Delta(H)}\varepsilon^{-2}\right).$
@@
Consider a motif $H$ on $k$ vertices. By Section 2.3, for $p=\Omega\left(n^{-1/m_{H}}\right)$ $\mathcal{G}_{n,p}$ satisfies the properties that all $H^{\prime}\subseteq H$ the number of labeled induced copies of $H^{\prime}$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{i}C_{H^{\prime}}$ for some $C_{H^{\prime}}$ with high constant probability, and that $r_{\ell-1}(H)\geq r_{\ell}(H)$ for all $\ell\in\{2,3,\dots,k\}$. Therefore, Theorem 5.1 implies that quality control for the count of $H$ in an input graph $G$ compared to $\mathcal{G}_{n,p}$ can be accomplished in $O\left(r_{k}(H)^{-2}\varepsilon^{-4}\right)=p^{-O(\Delta(H))}\cdot\mathrm{poly}(1/\varepsilon)$ queries. ∎
@@ 4.3 Algorithmic complexity
$p=\Omega\left(\max\left\{\left(\log(n)/n\right)^{1/(4\sigma(H)-1)},n^{-1/m_{H}^{\star}}\right\}\right)$, and induced motif $H$ on $k$ vertices, there is an algorithm for $(\mathcal{G}_{n,p},\rho_{H})$-quality control that makes $\frac{1}{p^{O(\Delta(H))}}$ adjacency matrix queries and has a runtime of $\frac{1}{p^{O(\Delta(H))}}$.
@@ 4.3.2 Moving to induced motifs
We also observe that Section 2.3 can imply that, for $p=\Omega\left(\max_{H^{\prime}\supseteq H,|H^{\prime}|=k}n^{-1/m_{H^{\prime}}}\cdot\varepsilon^{-2}\right)$, with probability at least $9/10$, for all $H^{\prime}\supseteq H,|H^{\prime}|=k$, the number of copies of $H^{\prime}$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100k^{k}}\binom{n}{k}p^{e(H^{\prime})}(1-p)^{\binom{k}{2}-e(H)}$. This is because, as noted in Section 2 after Section 2.3, the choice of $\varepsilon/100$ was not specific and can be changed with the hidden constants in the lower bound on $p$. The same applies for the error probability.
@@
Completeness: Suppose $G\sim\mathcal{G}_{n,p}$. Then, by the observation made above, with probability at least $9/10$, for all $H^{\prime}\supseteq H,|H^{\prime}|=k$, the number of copies of $H^{\prime}$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100k^{k}}\binom{n}{k}p^{e(H^{\prime})}(1-p)^{\binom{k}{2}-e(H)}$. Additionally, with similar (up to constants) parameter choices for the algorithm, the correctness of Algorithm NonInduced-Motif-Quality-Efficient on an input $H$ can be boosted to $1-1/(10\cdot 2^{\binom{k}{2}})$. Therefore, by the correctness of NonInduced-Motif-Quality-Efficient, with probability at least $9/10$, all calls to Algorithm NonInduced-Motif-Quality-Efficient in Step 1 of Algorithm Induced-Motif-Quality-Efficient accept. Therefore, Algorithm Induced-Motif-Quality-Efficient accepts $G\sim\mathcal{G}_{n,p}$ with high probability.
@@
Therefore, there exists an $H^{\prime}\supset H,|H^{\prime}|=k$ such that $\left|\widetilde{C}_{H^{\prime}}(G)-\widetilde{\mu}_{H^{\prime}}\right|\geq\varepsilon\mu_{H}/2^{\binom{k}{2}}$. We refer to this $H^{\prime}$ as $H^{\circ}$
@@ 5 Quality control for other random graph models
In this section, we generalize our results to the setting of quality control of any motif $H$ as compared to any distribution $\mathcal{D}_{n}$ over graphs with a good concentration of motif counts. Particularly, we will consider distributions $\mathcal{D}_{n}$ over graphs such that, for $G\sim\mathcal{D}_{n}$, for all $P\subseteq H$ the number of labeled induced copies of $P$ is in the range $(1\pm\frac{\varepsilon}{100})E_{\mathcal{D}_{n}}(P)$ for some $E_{\mathcal{D}_{n}}(P)$ with high constant probability.
@@
    For $G\sim\mathcal{D}_{n}$, for all $\ell\in[k]$ and all $P\subseteq H$ on $\ell$ vertices, the number of labeled induced copies of $P$ is in the range $(1\pm\frac{\varepsilon}{100})E_{\mathcal{D}_{n}}(P)$ with high constant probability.

### v2

@@ 1.1 Quality Control problems
For function $\rho:\mathcal{X}\to\mathbb{R}_{\geq 0}$, small constant $\varepsilon>0$, and a distribution $\mathcal{D}:\mathcal{X}\to\mathbb{R}_{\geq 0}$ such that $\mathbb{P}_{x\sim\mathcal{D}}\left[\left|\rho(x)-1\right|\leq\varepsilon\right]\geq 1-o(1)$, a $(\mathcal{D},\rho)$-quality control algorithm $A$ must satisfy the following:
@@
    $\mathbb{P}_{x\sim\mathcal{D};R}\left[A(x)=\textsc{Accept}\right]\geq 1-o(1)$, where the probability is taken over distribution $\mathcal{D}$ and the randomness $R$ used by algorithm $A$.
@@
Above, we use $o(1)$ to mean that a quantity approaches 0 as the runtime of the algorithm tends to infinity. (Formally, the quality control algorithm takes a error parameter $\tau$ as input, and for every $\tau>0$ there exists a constant $c$ such that the quality control algorithm with error parameter $\tau$ accepts $x\sim\mathcal{D}$ with probability at least $1-\tau$ and the runtime of the quality control algorithm with parameter $\tau$ is at most $c$ times larger than the runtime with error parameter set to some fixed constant, say, $1/3$.)
@@ 1.3 Comparison to related works
Given this proximity to the notions of average case complexity, one may anticipate seeing instances of quality control problems arising naturally in the field. Indeed, one notable class of such problems is “refutations of random 3SAT” problems. Here a seminal definition of Feige [Fei02] requires that a refutation algorithm for 3SAT problem on distribution $\mathcal{D}$ should only accept unsatisfiable 3SAT inputs, while accepting most inputs drawn from $\mathcal{D}$. This corresponds exactly to a quality control problem with the parameter $\rho_{\mathrm{SAT}}(\phi)$ being $1$ if a formula $\phi$ is unsatisfiable and 0 otherwise.
Additionally, the problem of certifying properties of a random object has appeared in various contexts throughout theoretical computer science. Notably, natural proofs [RR97] examine such a problem (within the context of ruling out avenues towards super-polynomial circuit lower bounds). Additionally, as noted by [Tre06], certifying properties of random instances is a challenge and a crucial step in explicit constructions from random objects. The problem of certifying properties of random objects in these various settings can be framed as a quality control problem.
@@ 2.3 Properties of Erdős–Rényi graphs
For $G\sim\mathcal{G}_{n,p}$, with probability $1-o_{1/p}(1)$, the number of labeled induced copies of $H$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{v(H)}v(H)!\cdot p^{e(H)}(1-p)^{\binom{v(H)}{2}-e(H)}$.
@@
For $G\sim\mathcal{G}_{n,p}$ and $p=\omega\left(n^{-2/(k-1)}\cdot\varepsilon^{-2}\right)$, the number of $k$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{k}k!\cdot p^{\binom{k}{2}}$ with probability $1-o_{1/p}(1)$.
@@
For all $p=\Omega\left(\frac{1}{n}\right)$, there exists an $N$ such that for all $n\geq N$, $\mathbb{P}_{G\sim\mathcal{G}_{n,p}}\left(\alpha(G)\leq 2\frac{e(G)}{n-1}+1\right)=1-o_{1/p}(1)$.
@@
$\mathbb{P}_{G\sim\mathcal{G}_{n,p}}\left(\alpha(G)\leq 2np\right)=1-o_{1/p}(1).$
@@
Let $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$ and $s\geq 1200C^{4}/p^{5k}$. Let $\delta=\frac{p^{2k}}{12C^{2}}$. With probability $1-o_{1/p}(1)$, for $G\sim\mathcal{G}_{n,p}$, at least a $(1-o_{1/p}(1))$-fraction of subgraphs $S$ of size $s$ satisfy the following:
@@
Therefore, the expected number of $S\subseteq[n],|S|=s$ such that $G_{S}$ is not good is at most $\frac{1}{s}\cdot\binom{n}{s}=o\left(\binom{n}{s}\right)$. Therefore, by Markov’s inequality, with probability at least $1-o_{1/p}(1)$, at least a $(1-o_{1/p}(1))$-fraction of subgraphs of size $s$ are good. ∎
@@
Let $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$, $s\geq 1200C^{4}/p^{5k}$, and $\delta=\frac{p^{2k}}{12C^{2}}$. Consider $G\sim\mathcal{G}_{n,p}$, and choose a random subgraph $G_{S}$ of $G$ on $s$ vertices. Then, with probability $1-o_{1/p}(1)$, the subgraph on these $s$ vertices satisfies the following:
@@
This corollary follows by taking a union bound over the events in Section 2.3 that $G\sim\mathcal{G}_{n,p}$ does not satisfy the property that a $(1-o_{1/p}(1))$-fraction of size-$s$ subgraphs are good and that a non-good subgraph is chosen at random. ∎
@@ 3 Quality control of k-clique count
    When $p$ is additionally $\Omega(\log(n)/n)^{1/(4k-1)}$, the problem is solvable in $O(p^{-c_{3}k})$ time.
@@ 3.1.2 Proof of Inductive Lemma given Concentration Lemma
Suppose that a graph $G$ satisfies the property that for all $\ell\in\{2,3,\dots,k\}$ the number of labeled $\ell$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{4(k+1)})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$. Then, for size $s_{\ell}$ multisets, for all $\ell\in[k-1]$, $G$ is $\left((\ell+1)\cdot\frac{\varepsilon}{4(k+1)},(\ell+1)^{2}\cdot\frac{\varepsilon^{2}}{(k+1)^{2}}\cdot\frac{p^{2\ell-2}}{2048},s_{\ell}\right)$-exponentially robustly quasirandom with respect to $\ell$-clique counts.
@@
Completeness. By Section 2.3, with high probability, for $G\sim\mathcal{G}_{n,p}$ with $p=\omega\left(\varepsilon^{-2}n^{-2/(k-1)}\right)$, for all $\ell\leq k$ the number of labeled $\ell$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{4(k+1)})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$. In the following, suppose that the high-probability event that the clique counts are all in these typical ranges holds. Then, by Claim 3.1.3, for all $\ell\in[k]$, $G$ is $\left(\delta_{\ell},\nu_{\ell},s_{\ell}\right)$-exponentially robustly quasirandom with respect to $\ell$-clique counts, for
@@
By Section 3.1.2, from this exponential quasirandomness, we can conclude that, with probability at least $1-k\cdot 4^{k}\exp\left(-\nu_{k}s_{*}\right)=1-o(1)$, for all $\ell\in[k]$, the uniformly sampled multiset $S$ satisfies $C_{\ell}(S)\in(1\pm\frac{\varepsilon}{4})E_{s,\ell}$. Therefore, the algorithm accepts $G\sim\mathcal{G}_{n,p}$ with high probability.
@@ 3.3 Algorithmic complexity
Completeness. We first argue that, with high probability, $G\sim\mathcal{G}_{n,p}$ is accepted by Algorithm Clique-Quality-Efficient. Consider $G\sim\mathcal{G}_{n,p}$, and choose a random subgraph of $G$ on $s_{*}$ vertices. By Corollary 2.3, for $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$ and $s_{*}\geq 1200C^{4}/p^{5k}$, with probability 1o(1)1-o(1), the subgraph on these $s_{*}$ vertices satisfies the conditions checked for in Step 3 with high probability, as proven in Section 2.3. Therefore, Step 3 does not reject $G\sim\mathcal{G}_{n,p}$ with high probability.
@@ 4.1 Query complexity upper bound
Define $s_{*}$ as in Section 5. Then $s_{*}=\widetilde{O}\left(r_{k}(H)^{-1}\varepsilon^{-2}\right)=\widetilde{O}\left(p^{-2\Delta(H)}\varepsilon^{-2}\right).$
@@
Consider a motif $H$ on $k$ vertices. By Section 2.3, for $p=\omega\left(n^{-1/m_{H}}\right)$ $\mathcal{G}_{n,p}$ satisfies the properties that all $H^{\prime}\subseteq H$ the number of labeled induced copies of $H^{\prime}$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{i}C_{H^{\prime}}$ for some $C_{H^{\prime}}$ with high probability, and that $r_{\ell-1}(H)\geq r_{\ell}(H)$ for all $\ell\in\{2,3,\dots,k\}$. Therefore, Theorem 5.1 implies that quality control for the count of $H$ in an input graph $G$ compared to $\mathcal{G}_{n,p}$ can be accomplished in $O\left(r_{k}(H)^{-2}\varepsilon^{-4}\right)=p^{-O(\Delta(H))}\cdot\mathrm{poly}(1/\varepsilon)$ queries. ∎
@@ 4.3 Algorithmic complexity
$p$ in both $\Omega\left(\left(\log(n)/n\right)^{1/(4\sigma(H)-1)}\right)$ and $\omega\left(n^{-1/m_{H}^{\star}}\right)$, and induced motif $H$ on $k$ vertices, there is an algorithm for $(\mathcal{G}_{n,p},\rho_{H})$-quality control that makes $\frac{1}{p^{O(\Delta(H))}}$ adjacency matrix queries and has a runtime of $\frac{1}{p^{O(\Delta(H))}}$.
@@ 4.3.2 Moving to induced motifs
We also observe that Section 2.3 can imply that, for $p=\omega\left(\max_{H^{\prime}\supseteq H,|H^{\prime}|=k}n^{-1/m_{H^{\prime}}}\cdot\varepsilon^{-2}\right)$, with probability at least 1o(1)1-o(1), for all $H^{\prime}\supseteq H,|H^{\prime}|=k$, the number of copies of $H^{\prime}$ in $G\sim\mathcal{G}_{n,p}$ is in the range $(1\pm\frac{\varepsilon}{100k^{k}})\binom{n}{k}p^{e(H^{\prime})}(1-p)^{\binom{k}{2}-e(H)}$. This is because, as noted in Section 2 after Section 2.3, the choice of $\varepsilon/100$ was not specific and can be changed with the hidden constants in the lower bound on $p$.
@@
Completeness: Suppose $G\sim\mathcal{G}_{n,p}$. Then, by the observation made above, with probability 1o(1)1-o(1), for all $H^{\prime}\supseteq H,|H^{\prime}|=k$, the number of copies of $H^{\prime}$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100k^{k}})\binom{n}{k}p^{e(H^{\prime})}(1-p)^{\binom{k}{2}-e(H)}$. Therefore, by the correctness of NonInduced-Motif-Quality-Efficient, with probability 1o(1)1-o(1), all calls to Algorithm NonInduced-Motif-Quality-Efficient in Step 1 of Algorithm Induced-Motif-Quality-Efficient accept. Therefore, Algorithm Induced-Motif-Quality-Efficient accepts $G\sim\mathcal{G}_{n,p}$ with high probability.
@@
Therefore, there exists an $H^{\prime}\supset H,|H^{\prime}|=k$ such that $\left|\widetilde{C}_{H^{\prime}}(G)-\widetilde{\mu}_{H^{\prime}}\right|\geq\varepsilon\mu_{H}/2^{\binom{k}{2}}$. We refer to this $H^{\prime}$ as $H^{\circ}$.
@@ 5 Quality control for other random graph models
In this section, we generalize our results to the setting of quality control of any motif $H$ as compared to any distribution $\mathcal{D}_{n}$ over graphs with a good concentration of motif counts. Particularly, we will consider distributions $\mathcal{D}_{n}$ over graphs such that, for $G\sim\mathcal{D}_{n}$, for all $P\subseteq H$ the number of labeled induced copies of $P$ is in the range $(1\pm\frac{\varepsilon}{100})E_{\mathcal{D}_{n}}(P)$ for some $E_{\mathcal{D}_{n}}(P)$ with high probability.
@@
    For $G\sim\mathcal{D}_{n}$, for all $\ell\in[k]$ and all $P\subseteq H$ on $\ell$ vertices, the number of labeled induced copies of $P$ is in the range $(1\pm\frac{\varepsilon}{100})E_{\mathcal{D}_{n}}(P)$ with probability 1o(1)1-o(1).